# OpenOOD v1.5: Enhanced Benchmark for Out-of-Distribution Detection

### v4

@@ 1.1 Background and Motivation
2 Problem Statement and Evaluation Protocol
This work considers OOD detection in the context of (supervised) multi-class image classification. In this section, we formally describe the problem statement and present the evaluation protocol specified by OpenOOD for both standard and full-spectrum setting.
2.1 Standard OOD Detection
There are two goals in standard OOD detection. We first establish context. In each image classification problem, there exists a pre-defined set of semantic categories that we expect the model to identify (e.g., dog and automobile in CIFAR-10), which defines what images are considered in-distribution (ID). We refer to the group of labels and the associated joint distribution as $\mathcal{Y}_{\text{ID}}$ and $\mathcal{D}_{\text{ID}}$, respectively, where $\forall(\mathbf{x},y)\sim\mathcal{D}_{\text{ID}},y\in\mathcal{Y}_{\text{ID}}$. In an open world, there exists semantic groups that don’t belong to the pre-defined and finite $\mathcal{Y}_{\text{ID}}$, which forms the out-of-distribution (OOD) space $\mathcal{Y}_{\text{OOD}}=\{y|y\notin\mathcal{Y}_{\text{ID}}\}$ and $\mathcal{D}_{\text{OOD}}$.
Goal 1 The first goal of OOD detection is to have a discriminative model that accurately classifies ID samples drawn from $\mathcal{D}_{\text{ID}}$. To evaluate this goal, we use widely adopted image classification datasets (such as CIFAR and ImageNet) as the target distribution $\mathcal{D}_{\text{ID}}$. As in normal classification tasks, we train the classifier with specific methods by minimizing the empirical risk on the training dataset $\mathcal{D}_{\text{ID}}^{\texttt{train}}$ and evaluate its performance on the unseen ID test images $\mathcal{D}_{\text{ID}}^{\texttt{test}}$. The metric for goal 1 is classification accuracy.
Goal 2 The second goal is to have a detector module (typically built upon the trained classifier) that accurately identifies whether an incoming image in the inference phase is ID or OOD, i.e., $y\in\mathcal{Y}_{\text{ID}}$ or $y\notin\mathcal{Y}_{\text{ID}}$. Following common practices, we take publicly available datasets whose categories are OOD w.r.t. $\mathcal{Y}_{\text{ID}}$ as the source of OOD samples $\mathcal{D}_{\text{OOD}}$. Since certain methods incorporate OOD images at training time to improve OOD detection capability [27, 28, 29, 30], we make disctinction between OOD training samples $\mathcal{D}_{\text{OOD}}^{\texttt{train}}$ and OOD test samples $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$, where they should have non-overlapping categories between each other (i.e., $\mathcal{Y}_{\text{OOD}}^{\texttt{train}}\cap\mathcal{Y}_{\text{OOD}}^{\texttt{test}}=\varnothing$) to avoid trivial evaluation [27]. For each ID test set $\mathcal{D}_{\text{ID}}^{\texttt{test}}$, we use multiple OOD test sets $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$ and measure the OOD detection performance on each ID-OOD pair. We treat OOD samples as positive and ID samples as negative. We use the well established 1) area under the receiver operating characteristic (AUROC), 2) area under the Precision-Recall curve (AUPR), and 3) false positive rate at 95% true negative rate (FPR@95) as metrics for goal 2. AUROC and AUPR are threshold-independent measurements, while FPR95 reflects the performance at a specific threshold.
Remark. There are two extra efforts that make our evaluation protocol much more rigorous than prior works. First, OpenOOD divides considered OOD datasets into two groups: near-OOD and far-OOD (or equivalently, hard-OOD and easy-OOD). The grouping is based on either image semantics or empirical difficulty, which can give a more comprehensive evaluation of OOD detectors in the face of different types of OOD samples (see Appendix A for details). Second, we introduce ID validation samples $\mathcal{D}_{\text{ID}}^{\texttt{val}}$ and OOD validation samples $\mathcal{D}_{\text{OOD}}^{\texttt{val}}$ in all evaluation settings. This step has been long overlooked by existing works as they often tune hyperparameters or select models directly using ID test samples $\mathcal{D}_{\text{ID}}^{\texttt{test}}$ and OOD test samples $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$ [5, 31, 32, 6], which goes against basic machine learning principles and leads to unfair and overoptimistic results. In contrast, we use $\mathcal{D}_{\text{ID}}^{\texttt{val}}$ and $\mathcal{D}_{\text{OOD}}^{\texttt{val}}$ whenever applicable to ensure fairness, where $\mathcal{D}_{\text{ID}}^{\texttt{val}}$ is a small subset held out from $\mathcal{D}_{\text{ID}}^{\texttt{test}}$, and $\mathcal{D}_{\text{OOD}}^{\texttt{val}}$ is carefully constructed such that $\mathcal{Y}_{\text{OOD}}^{\texttt{val}}\cap\mathcal{Y}_{\text{OOD}}^{\texttt{test}}=\varnothing$.
Figure 3: Illustration of full-spectrum OOD detection [29]. Standard detection only concerns semantic shift by detecting (c) + (d) from (a), while full-spectrum detection takes into account covariate shift and aims to separate (c) + (d) from (a) + (b). An ideal system should be robust to the non-semantic covariate shift (OOD generalization) while being able to identify semantic shift (OOD detection).
2.2 Full-Spectrum OOD Detection
Unlike standard OOD detection which only concerns semantic shift between the training and test distribution, full-spectrum detection [19] further takes into account the non-semantic covariate shift by considering covariate-shifted ID (csID) images [20, 33, 21] (see Figure 3). Covariate-shifted samples $\mathbf{x}_{\text{csID}}$ are non-i.i.d. w.r.t. ID samples $\mathbf{x}_{\text{ID}}$, yet they still semantically belong to one of the ID categories, i.e., $y_{\text{csID}}\in\mathcal{Y}_{\text{ID}}$. Concrete examples of csID images include the ones that exhibit slight corruptions [20] and style changes [21].
In the full-spectrum setting, goal 1 of OOD detection becomes accurate classification of both ID and csID samples from $\mathcal{D}_{\text{ID}}^{\texttt{test}}$ and $\mathcal{D}_{\text{csID}}^{\texttt{test}}$. Meanwhile, goal 2 turns into that the OOD detector should be able to flag samples from $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$ as OOD, while identifying samples from $\mathcal{D}_{\text{ID}}^{\texttt{test}}$ and $\mathcal{D}_{\text{csID}}^{\texttt{test}}$ as ID. We leverage established covariate-shifted datasets such as ImageNet-C [20] and ImageNet-R [21] to serve as $\mathcal{D}_{\text{csID}}^{\texttt{test}}$. The metrics for both goals remain the same as in standard OOD detection.
Remark. Essentially, the full-spectrum setting puts OOD detection into a broader scope by simultaneously considering two important distributional shifts. The standard detection setting (adopted by most prior works) ignores the covariate shift that might be encountered in the inference, and it remains unclear how current OOD detectors will respond to covariate-shifted samples. Our later results show that nearly all methods face significant performance degradation in full-spectrum detection, and we believe that this is an important problem to address.
3 Supported Benchmarks and Methods
In this section, we provide an overview of the supported benchmarks and methods in OpenOOD v1.5. Due to space limit, detailed descriptions of each benchmark and method are left in Appendix A and Appendix B, respectively.
Benchmarks. OpenOOD v1.5 supports 6 benchmarks, including 4 for standard OOD detection and 2 for full-spectrum OOD detection. The standard benchmarks are constructed with CIFAR-10, CIFAR-100, ImageNet-200, and ImageNet-1K being the ID data, respectively. The full-spectrum benchmarks are built around ImageNet-200 and ImageNet-1K, respectively. Among them, CIFAR and ImageNet-1K standard benchmark are adapted from v1 release with necessary changes for fairness and usefulness. ImageNet-200 and full-spectrum benchmarks are newly introduced in v1.5.
Comparison with prior OOD and OSR benchmarks. Most of the OOD datasets considered in previous works fall into our far-OOD category, meaning that the more difficult near-OOD detection was less emphasized than our benchmarks do. Meanwhile, we exclude problematic OOD datasets (e.g., the resized version of LSUN and TIN [5] with obvious artifacts [7]) which make detection trivial and much less meaningful [7]. Following one of the seminal works [34], OSR papers often construct ID-OOD pairs by partitioning a single dataset into two splits (e.g., using a 6-class subset of CIFAR-10 as ID and the other 4-class subset as OOD). While such practice is well-suited for studying near-OOD detection [35], it inevitably reduces the scale and complexity of the problem, making the resulted benchmarks less representative for real-world scenarios. In contrast, our benchmarks (especially ImageNet ones) operate at a much larger scale (in terms of the number of ID and OOD categories) than conventional OSR benchmarks.
Supported methods. By including more recent works, OpenOOD v1.5 now implements in total 40 advanced OOD detection methods within our framework. They are categorized into four groups. Post-hoc inference methods designs post-processors that are applied to the base classifier to generate the “OOD score” (which is then thresholded to produce the binary ID/OOD prediction). They only take effect at inference phase and by default assume that the classifier is trained with the standard cross-entropy loss. In contrast, training methods involve training-time regularization. Most of them assume no access to the auxiliary OOD training data (w/o outlier data), while some methods do (w/ outlier data). We also consider several data augmentation methods.
4 Experiment Setup
Training. For CIFAR-10/100 and ImageNet-200, we train a ResNet-18 [36] for 100 epochs. We consider the standard cross-entropy training for post-hoc methods . The optimizer is SGD with a momentum of 0.9. We use a learning rate of 0.1 with cosine annealing decay schedule [37]. A weight decay of 0.0005 is applied. The batch size is 128 for CIFAR-10/100 and 256 for ImageNet-200. Some methods have specific setup, and we adopt their official implementations and hyperparameters whenever possible.
For ImageNet-1K, we evaluate post-hoc methods with pre-trained models from torchvision [38]. In addition to ResNet-50 that is considered in OpenOOD v1, v1.5 further includes ViT [39] and Swin Transformer [40] for comprehensive evaluation. For training methods, we focus on ResNet-50 and use official checkpoints when possible. Otherwise, we fine-tune the torchvision pre-trained checkpoint for 30 epochs with a learning rate of 0.001. Again, we use a batch size of 256 and weight decay of 0.0005. CIFAR and ImageNet models are trained using 1 and 2 Quadro RTX 6000 GPUs (24GB memory), respectively. Except ImageNet-1K experiments, we perform 3 independent training runs for each method. All training scripts are thoroughly documented in our Github repo.
Evaluation. As aforementioned, we use AUROC, AUPR, and FPR@95 as metrics. In the paper we focus on near- and far-OOD AUROC which are averaged over all OOD datasets in each group. AUROC can be interpreted as the probability that the detector correctly separates ID and OOD samples; the random-guessing baseline is 50%, and the higher the better. Results under other metrics and per-dataset statistics are available in our full result table. For post-hoc methods, OpenOOD supports automatic hyperparameter search using ID and OOD validation samples. The hyperparameter that yields the best AUROC is used for the final test. Evaluation scripts can be found here.
Notes on missing results. The main results for standard OOD detection are presented in Table 1. A few numbers are missing for the following reasons. OpenGAN [6] has not shown success on ImageNet-1K, and substantial changes are required to make it work with ImageNet models. CSI [7], VOS [41], and NPOS [42] are infeasible with our compute resources on ImageNet. CIDER [43] and NPOS trains the CNN backbone without the final linear classifier, and the exact code for evaluating ID accuracy is not provided in their official implementations. Lastly, we do not consider training with outlier data $\mathcal{D}_{\text{OOD}}^{\texttt{train}}$ on ImageNet-1K since it is difficult to find OOD training samples that do not overlap with test OOD data, and no relevant methods have considered ImageNet-1K.
Table 1: Main results from OpenOOD v1.5 on standard OOD detection. In this table we use AUROC as the metric for OOD detection. Whenever applicable, we report the average number and the corresponding standard deviation obtained from 3 training runs. The best result within each group is bolded. Results for data augmentation methods are listed in Table 2. Full result table including other metrics and per-dataset statistics can be found here.
CIFAR-10 CIFAR-100 ImageNet-200 ImageNet-1K
- Post-hoc Inference Methods
OpenMax [2] (CVPR’16) 87.62(±0.29) 89.62(±0.19) 95.06(±0.30) 76.41(±0.25) 79.48(±0.41) 77.25(±0.10) 80.27(±0.10) 90.20(±0.17) 86.37(±0.08) 74.77 89.26 76.18
MSP [1] (ICLR’17) 88.03(±0.25) 90.73(±0.43) 95.06(±0.30) 80.27(±0.11) 77.76(±0.44) 77.25(±0.10) 83.34(±0.06) 90.13(±0.09) 86.37(±0.08) 76.02 85.23 76.18
TempScale [44] (ICML’17) 88.09(±0.31) 90.97(±0.52) 95.06(±0.30) 80.90(±0.07) 78.74(±0.51) 77.25(±0.10) 83.69(±0.04) 90.82(±0.09) 86.37(±0.08) 77.14 87.56 76.18
ODIN [5] (ICLR’18) 82.87(±1.85) 87.96(±0.61) 95.06(±0.30) 79.90(±0.11) 79.28(±0.21) 77.25(±0.10) 80.27(±0.08) 91.71(±0.19) 86.37(±0.08) 74.75 89.47 76.18
MDS [31] (NeurIPS’18) 84.20(±2.40) 89.72(±1.36) 95.06(±0.30) 58.69(±0.09) 69.39(±1.39) 77.25(±0.10) 61.93(±0.51) 74.72(±0.26) 86.37(±0.08) 55.44 74.25 76.18
MDSEns [31] (NeurIPS’18) 60.43(±0.26) 73.90(±0.27) 95.06(±0.30) 46.31(±0.24) 66.00(±0.69) 77.25(±0.10) 54.32(±0.24) 69.27(±0.57) 86.37(±0.08) 49.67 67.52 76.18
RMDS [45] (arXiv’21) 89.80(±0.28) 92.20(±0.21) 95.06(±0.30) 80.15(±0.11) 82.92(±0.42) 77.25(±0.10) 82.57(±0.25) 88.06(±0.34) 86.37(±0.08) 76.99 86.38 76.18
Gram [46] (ICML’20) 58.66(±4.83) 71.73(±3.20) 95.06(±0.30) 51.66(±0.77) 73.36(±1.08) 77.25(±0.10) 67.67(±1.07) 71.19(±0.24) 86.37(±0.08) 61.70 79.71 76.18
EBO [47] (NeurIPS’20) 87.58(±0.46) 91.21(±0.92) 95.06(±0.30) 80.91(±0.08) 79.77(±0.61) 77.25(±0.10) 82.50(±0.05) 90.86(±0.21) 86.37(±0.08) 75.89 89.47 76.18
OpenGAN [6] (ICCV’21) 53.71(±7.68) 54.61(±15.51) 95.06(±0.30) 65.98(±1.26) 67.88(±7.16) 77.25(±0.10) 59.79(±3.39) 73.15(±4.07) 86.37(±0.08) N/A N/A N/A
GradNorm [48] (NeurIPS’21) 54.90(±0.98) 57.55(±3.22) 95.06(±0.30) 70.13(±0.47) 69.14(±1.05) 77.25(±0.10) 72.75(±0.48) 84.26(±0.87) 86.37(±0.08) 72.96 90.25 76.18
ReAct [13] (NeurIPS’21) 87.11(±0.61) 90.42(±1.41) 95.06(±0.30) 80.77(±0.05) 80.39(±0.49) 77.25(±0.10) 81.87(±0.98) 92.31(±0.56) 86.37(±0.08) 77.38 93.67 76.18
MLS [14] (ICML’22) 87.52(±0.47) 91.10(±0.89) 95.06(±0.30) 81.05(±0.07) 79.67(±0.57) 77.25(±0.10) 82.90(±0.04) 91.11(±0.19) 86.37(±0.08) 76.46 89.57 76.18
KLM [14] (ICML’22) 79.19(±0.80) 82.68(±0.21) 95.06(±0.30) 76.56(±0.25) 76.24(±0.52) 77.25(±0.10) 80.76(±0.08) 88.53(±0.11) 86.37(±0.08) 76.64 87.60 76.18
VIM [15] (CVPR’22) 88.68(±0.28) 93.48(±0.24) 95.06(±0.30) 74.98(±0.13) 81.70(±0.62) 77.25(±0.10) 78.68(±0.24) 91.26(±0.19) 86.37(±0.08) 72.08 92.68 76.18
KNN [16] (ICML’22) 90.64(±0.20) 92.96(±0.14) 95.06(±0.30) 80.18(±0.15) 82.40(±0.17) 77.25(±0.10) 81.57(±0.17) 93.16(±0.22) 86.37(±0.08) 71.10 90.18 76.18
DICE [49] (ECCV’22) 78.34(±0.79) 84.23(±1.89) 95.06(±0.30) 79.38(±0.23) 80.01(±0.18) 77.25(±0.10) 81.78(±0.14) 90.80(±0.31) 86.37(±0.08) 73.07 90.95 76.18
RankFeat [50] (NeurIPS’22) 79.46(±2.52) 75.87(±5.06) 95.06(±0.30) 61.88(±1.28) 67.10(±1.42) 77.25(±0.10) 56.92(±1.59) 38.22(±3.85) 86.37(±0.08) 50.99 53.93 76.18
ASH [17] (ICLR’23) 75.27(±1.04) 78.49(±2.58) 95.06(±0.30) 78.20(±0.15) 80.58(±0.66) 77.25(±0.10) 82.38(±0.19) 93.90(±0.27) 86.37(±0.08) 78.17 95.74 76.18
SHE [18] (ICLR’23) 81.54(±0.51) 85.32(±1.43) 95.06(±0.30) 78.95(±0.18) 76.92(±1.16) 77.25(±0.10) 80.18(±0.25) 89.81(±0.61) 86.37(±0.08) 73.78 90.92 76.18
- Training Methods (w/o Outlier Data)
ConfBranch [51] (arXiv’18) 89.84(±0.24) 92.85(±0.29) 94.88(±0.05) 71.60(±0.62) 68.90(±1.83) 76.59(±0.27) 79.10(±0.24) 90.43(±0.18) 85.92(±0.07) 70.66 83.94 75.63
RotPred [52] (NeurIPS’19) 92.68(±0.27) 96.62(±0.18) 95.35(±0.52) 76.43(±0.16) 88.40(±0.13) 76.03(±0.38) 81.59(±0.20) 92.56(±0.09) 86.37(±0.16) 76.52 90.00 76.55
G-ODIN [32] (CVPR’20) 89.12(±0.57) 95.51(±0.31) 94.70(±0.25) 77.15(±0.28) 85.67(±1.58) 74.46(±0.04) 77.28(±0.10) 92.33(±0.11) 84.56(±0.28) 70.77 85.51 74.85
CSI [7] (NeurIPS’20) 89.51(±0.19) 92.00(±0.30) 91.16(±0.14) 71.45(±0.27) 66.31(±1.21) 61.60(±0.46) N/A N/A N/A N/A N/A N/A
ARPL [53] (TPAMI’21) 87.44(±0.15) 89.31(±0.32) 93.66(±0.11) 74.94(±0.93) 73.69(±1.80) 70.70(±1.08) 82.02(±0.10) 89.23(±0.11) 83.95(±0.32) 76.30 85.50 75.87
MOS [54] (CVPR’21) 71.45(±3.09) 76.41(±5.93) 94.83(±0.37) 80.40(±0.18) 80.17(±1.21) 76.98(±0.20) 69.84(±0.46) 80.46(±0.92) 85.60(±0.20) 72.85 82.75 72.81
VOS [41] (ICLR’22) 87.70(±0.48) 90.83(±0.92) 94.31(±0.64) 80.93(±0.29) 81.32(±0.09) 77.20(±0.10) 82.51(±0.11) 91.00(±0.28) 86.23(±0.19) N/A N/A N/A
LogitNorm [55] (ICML’22) 92.33(±0.08) 96.74(±0.06) 94.30(±0.25) 78.47(±0.31) 81.53(±1.26) 76.34(±0.17) 82.66(±0.15) 93.04(±0.21) 86.04(±0.15) 74.62 91.54 76.45
CIDER [43] (ICLR’23) 90.71(±0.16) 94.71(±0.36) N/A 73.10(±0.39) 80.49(±0.68) N/A 80.58(±1.75) 90.66(±1.68) N/A 68.97 92.18 N/A
NPOS [42] (ICLR’23) 89.78(±0.33) 94.07(±0.49) N/A 78.35(±0.37) 82.29(±1.55) N/A 79.40(±0.39) 94.49(±0.07) N/A N/A N/A N/A
- Training Methods (w/ Outlier Data)
OE [27] (NeurIPS’18) 94.82(±0.21) 96.00(±0.13) 94.63(±0.26) 88.30(±0.10) 81.41(±1.49) 76.84(±0.42) 84.84(±0.16) 89.02(±0.18) 85.82(±0.21) N/A N/A N/A
MCD [28] (ICCV’19) 91.03(±0.12) 91.00(±1.10) 94.95(±0.04) 77.07(±0.32) 74.72(±0.78) 75.83(±0.04) 83.62(±0.09) 88.94(±0.10) 86.12(±0.17) N/A N/A N/A
UDG [29] (ICCV’21) 89.91(±0.25) 94.06(±0.90) 92.36(±0.84) 78.02(±0.10) 79.59(±1.77) 71.54(±0.64) 74.30(±1.63) 82.09(±2.78) 68.11(±1.24) N/A N/A N/A
MixOE [30] (WACV’23) 88.73(±0.82) 91.93(±0.69) 94.55(±0.32) 80.95(±0.20) 76.40(±1.44) 75.13(±0.06) 82.62(±0.03) 88.27(±0.41) 85.71(±0.07) N/A N/A N/A
ImageNet-200 ImageNet-1K
MSP [1] ASH [17] ID Acc. MSP [1] ReAct [13] ASH [17] ID Acc.
StyleAugment [22] 80.99 / 88.44 80.65 / 93.70 83.41 75.78 / 85.73 76.70 / 91.88 78.21 / 94.90 74.68
RandAugment [23] 83.17 / 90.34 81.56 / 94.53 86.58 76.60 / 85.27 78.30 / 93.50 79.81 / 95.01 76.90
AugMix [24] 83.49 / 90.68 82.87 / 94.66 87.01 77.49 / 86.67 79.94 / 93.70 82.16 / 96.05 77.63
DeepAugment [21] 81.39 / 88.79 80.61 / 93.84 85.00 76.67 / 86.26 78.43 / 92.12 79.14 / 93.90 76.77
PixMix [25] 82.15 / 90.23 81.36 / 95.01 85.79 76.86 / 85.63 79.12 / 91.59 78.92 / 92.17 77.44
RegMixup [26] 84.13 / 90.81 79.38 / 92.74 87.25 77.04 / 86.31 77.68 / 92.45 78.45 / 95.35 76.68
5 Analysis
In this section, we discuss multiple observations and insights that arise from the comprehensive results provided by OpenOOD v1.5. We start with standard OOD detection before delving into the full-spectrum setting.
5.1 Standard Detection
No single winner. In Table 1, there is no single method that consistently outperforms others across benchmarks, and the ranking can be quite different from one dataset to another. For example, ReAct [13] and ASH [17] are extremely powerful on ImageNet but less competitive on CIFAR. In contrast, methods that yield remarkable performance on small datasets (e.g., KNN [16] and RotPred [52]) do not show clear advantage on large datasets. This observation further emphasizes the importance of evaluation in large-scale settings as they better approximate complex scenarios in the real world.
Data augmentations help. While data augmentations have been shown beneficial for standard classification [23] and OOD generalization [22, 24, 21, 25, 26], their effects for OOD detection remain unclear. In Table 2, we find that several data augmentation methods, despite not being designed to improve OOD detection, can actually boost detection rates in many cases. More interestingly, the performance gain is amplified when they are combined with powerful post-processors. For example, compared with the baseline of cross-entropy training on ImageNet-1K near-OOD, AugMix [24] achieves 76.02 + 1.47 = 77.49% AUROC and 78.17 + 3.99 = 82.16% AUROC when working with MSP [1] and ASH [17], respectively. 82.16% is the current best score among all methods. The results indicate that the effects from data augmentations and post-processors are complementary.
Vision transformers do not outperform ResNets. We visualize in Figure 5 the performance of a few powerful post-hoc methods on ImageNet-1K with ResNet-50, ViT-B-16 [39], and Swin-T [40] as the classifier. We use the ImageNet-1K-pre-trained checkpoints provided by torchvision. As reference, they have 25.6M, 86.6M, and 28.3M parameters, and their ID accuracy is 76.18%, 81.14%, and 81.59%, respectively. Interestingly, despite their better ID classification performance, we find that vision transformers do not show noticeable improvements over ResNets for OOD detection, which aligns with the observation in [14]. Meanwhile, different post-processor may favor different architecture. For instance, the top-2 post-processor on ResNet-50, i.e., ASH [17] and ReAct [13], both have significant performance when operating on transformer. RMDS [45], in contrast, suits transformers much better than ResNets.
Training methods excel at small datasets. On CIFAR-10/100, we find that training-time regularizations (the second group in Table 1) can provide better OOD detection capability than post-hoc methods. In particular, RotPred [52] and LogitNorm [55] stand out as two powerful training methods (without using outlier data). On CIFAR-10, they both lead to $\sim$2and $\sim$3increase in near- and far-OOD AUROC, respectively, compared to the best-performing post-processors. Meanwhile, however, training methods in general do not outperform post-hoc ones on ImageNet-200/1K. This might be due to that more sophisticated training dynamics require larger models or longer training, and the exact reason needs further investigation in future works.
Outlier data helps in certain cases. Compared with methods that do not have such consideration, incorporating OOD training data (the last group in Table 1) is helpful mainly when the test OOD samples are similar to the training ones. For instance, OE [27] yields the highest near-OOD AUROC on CIFAR-100 because the used OOD training set (TIN-597; see Appendix A for details) is similar to one of the near-OOD test sets (TIN). In contrast, it does not seem to be beneficial for detecting far-OOD samples (which are quite different from TIN images) and actually underperforms several other methods which do not use the outlier data.
5.2 Full-Spectrum Detection
Full-spectrum detection poses challenge for current detectors. This can be seen in Figure 6, where the near-OOD AUROC of most methods decreases by >10% on ImageNet-1K (similar trend holds for far-OOD; see full results here). The performance drop suggests that existing OOD detectors can be sensitive to the non-semantic covariate shift and are likely to flag covariate-shifted ID samples as OOD. Such behaviour is not ideal because: 1) It does not align with human perception/decision (e.g., a human annotator classifying dog and car wouldn’t mark a covariate-shifted dog image as something unknown or novel). 2) It can harm the classifier’s generalization capability on covariate-shifted ID data, as often times it is assumed that the classifier would refrain from making predictions when the sample is identified as OOD (or at least the OOD flag associated with the prediction would indicate its unreliability). As a result, we firmly believe that full-spectrum detection is an important open problem for ensuring human-model alignment and the model’s practical reliability. One method that stands out in this resort is MOS [54], which utilizes a two-level semantic hierarchy to facilitate classification and OOD detection. MOS exhibits the smallest performance drop when changing from standard to full-spectrum detection (from 72.85% to 66.17% in AUROC) and serves as a strong baseline for future works to tackle full-spectrum detection.
Table 3: ImageNet-1K full-spectrum detection results of data augmentation methods. The numbers in each cell represent the near-OOD / far-OOD AUROC. Again, they are helpful especially when combined with approriate post-processors and when performing near-OOD detection.
MSP [1] GradNorm [56] SHE [18] ID Acc.
CrossEntropy 60.79 / 72.32 62.70 / 83.49 61.21 / 83.04 54.35
StyleAugment [22] 62.09 / 74.37 65.27 / 81.62 66.64 / 82.64 55.44
RandAugment [23] 61.36 / 72.07 63.27 / 76.08 64.41 / 76.68 55.57
AugMix [24] 63.14 / 74.62 67.10 / 81.29 69.66 / 83.06 57.46
DeepAugment [21] 63.51 / 75.40 65.66 / 76.27 68.27 / 78.85 57.82
PixMix [25] 62.51 / 73.47 61.07 / 70.00 65.02 / 77.03 57.27
RegMixup [26] 61.32 / 72.87 61.86 / 79.98 64.71 / 81.23 55.55
Data augmentations continue to help in full-spectrum settings. Similar to our observations in standard OOD detection, in Table 3 we see that data augmentations also boost full-spectrum detection rates especially when combined with powerful post-hoc methods. For example, compared to the cross-entropy baseline, while AugMix [24] increases near-OOD AUROC “only” by 2.35% with the MSP detector [1], it leads to a much significant improvement of 8.45% when working with SHE [18]. AugMix + SHE is the current best approach in terms of full-spectrum near-OOD AUROC on ImageNet-1K. In the meantime, we do notice that data augmentations do not clearly benefits full-spectrum far-OOD AUROC, and the reasons require future study. That said, our results demonstrate that in general “data augmentation + post-processor” is promising for both standard and full-spectrum OOD detection.
6 Conclusion and Discussion
In this work we present OpenOOD v1.5, which enhances its earlier version by 1) providing results on the large-scale ImageNet, 2) investigating full-spectrum detection, and 3) introducing new features such as the leaderboard. We highlight several insights regarding the current state and future directions for OOD detection: 1) A single winner that leads to superior performance across multiple benchmarks is still missing. 2) Data augmentations are quite effective for improving OOD detection capability, and the combination of data augmentation with strong post-processor seems promising. 3) Near-OOD remains more challenging than far-OOD, yet their improvement is proportional to each other. 4) Full-spectrum detection poses challenge for existing methods. We hope that the codebase, benchmark, evaluation results, and insights of OpenOOD can accelerate the progress and foster collective efforts towards advancing the state-of-the-art in OOD detection.
Related work. To the best of our knowledge, OpenOOD v1.5 is the only work that comprehensively evaluates a wide range of OOD detection methods on multiple benchmarks of various sizes. We discuss a few works that relate to OpenOOD in certain aspects in Appendix C.
Limitation. In this work we focus on the context where there assumes to be a discriminative classifier for the ID classification in the first place. As a result, all the OOD detection methods considered in our work are discriminative. OOD detection approaches that are based on generative modeling (e.g., [57, 58, 59, 60]) are not currently included in OpenOOD.
Societal Impact. OOD detection is an important topic for machine learning safety as it studies how deep neural networks can handle unknown inputs desirably. As an open-sourced, unified, and comprehensive benchmark for OOD detection, OpenOOD is expected to benefit the whole community and facilitate relevant research, which we believe has positive societal impacts.
In addition to the maintenance, in the future v2 release we plan to extend the scope of OpenOOD beyond image classification and include more application scenarios such as object detection, semantic segmentation, and natural language processing tasks. Specifically, it would be interesting to see whether current OOD detectors, which are designed for image classifiers, can generalize to different problems and modalities.
Acknowledgments and Disclosure of Funding
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
[…]

### v5

@@ Appendix A Evaluation Pitfalls of OOD Detection
Appendix B New Features and Updates of OpenOOD v1.5
OpenOOD v1.5 introduces new features including a leaderboard hosted online to track the-state-of-the-art based on various methods and a lightweight evaluator which enables easy evaluation with a few lines of code (see Figure 8). Other updates such as adding newer methods and fixing implementation bugs are documented in the changelog.101010 https://github.com/Jingkang50/OpenOOD/wiki/OpenOOD-v1.5-change-log
Appendix C Supported Methods
We now overview the supported methods of OpenOOD v1.5.
Post-Hoc Inference Methods. Recall from Equation 1 that given an input image, OOD detectors function by assigning an “OOD score” that is computed based on certain outputs from the base classifier, which will then be thresholded to give the binary prediction. The first line of works focus on designing such post-processors/scoring mechanisms that best separate ID and OOD samples. MSP Hendrycks and Gimpel (2017) takes the maximum softmax probability over ID categories as the score. OpenMax Bendale and Boult (2016) is a replacement for the softmax layer which directly estimates the probability of an input being from an unknown class. TempScale Guo et al. (2017) calibrates softmax probabilities with temperature scaling. ODIN Liang et al. (2018) further introduces input preprocessing on top of TempScale. MDS Lee et al. (2018) fits class-conditional Gaussian distribution on the penultimate layer features of the classifier and derives OOD score with Mahalanobis distance. MDSEns Lee et al. (2018) is another version of MDS which leverages multiple intermediate layers and forms a feature ensemble. RMDS Ren et al. (2021) improves MDS by considering the “background score” computed from an unconditional Gaussian distribution. Gram Sastry and Oore (2020) identifies abnormal patterns from the Gram Matrices of intermediate feature maps. EBO Liu et al. (2020) applies energy function to the logits to compute OOD score. OpenGAN Kong and Ramanan (2021) trains a GAN in the classifier’s feature space and uses the discriminator as the post-processor. GradNorm Huang et al. (2021) computes the KL divergence between the softmax probability distribution and the uniform distribution and takes the gradients of penultimate layer weights w.r.t. the KL divergence as OOD score. ReAct Sun et al. (2021) rectifies feature vectors by thresholding their elements with a certain magnitude. MLS Hendrycks et al. (2022) uses the maximum logit. KLM Hendrycks et al. (2022) looks at the KL divergence between the softmax probability distribution and a “template” distribution. VIM Wang et al. (2022) augments the logits with the norm of feature residual compared with ID training samples’ features to compute the OOD score. KNN Sun et al. (2022) applies KNN to the penultimate layer’s features. DICE Sun and Li (2022) sparsifies the last linear layer before computing the logits. RankFeat Song et al. (2022) transforms the feature matrices such that their rank is 1. ASH Djurisic et al. (2023) shapes later layer activations by removing a large portion of the elements and simplifing the rest. SHE Zhang et al. (2023b) maintains a template representation for each ID category and detects OOD samples by measuring the distance between the representation of an input to that template.
Training methods without outlier data. Unlike post-hoc methods that only interfere with the inference process, training methods involve training-time regularization to enhance OOD detection capability. ConfBranch DeVries and Taylor (2018) trains another branch in addition to the classification one to explicitly learn the estimate of model uncertainty. RotPred Hendrycks et al. (2019b) includes an extra head to predict the rotation angle of rotated inputs in a self-supervised manner, and the rotation head together with the classification head is used for OOD detection. G-ODIN Hsu et al. (2020) utilizes a dividend/divisor structure and decomposes the softmax confidence for better ID-OOD separation. CSI Tack et al. (2020) explores self-supervised contrastive learning objectives for OOD detectors. ARPL Chen et al. (2021) introduces “reciprocal points” for each ID category and trains the model by pushing the reciprocal point away from the corresponding ID cluster and encouraging OOD samples to gather around the reciprocal point. MOS Huang and Li (2021) incorporates a two-level hierarchical classifier and designs an accompanying OOD score to benefit OOD detection especially in large-scale settings. VOS Du et al. (2022) regularizes the feature space of the classifier under the assumption that the learned representations follow conditional Gaussian distributions. LogitNorm Wei et al. (2022) mitigates the over-confidence issue by training and testing with normalized logit vectors. CIDER Ming et al. (2023) regularizes the model’s hyperspherical space by increasing inter-class separability and intra-class compactness. NPOS Tao et al. (2023) is a non-parametric version of VOS which removes the Gaussian assumption and instead adopts KNN to model the feature distribution.
[…]